OpenStudy (anonymous):

Choose the equation that models the problem. Luisa won 17 tennis matches last summer. She played a total of 26 matches. How many matches did she lose overall? Let l represent the number of games that Luisa lost. A. l – 17 = 26 B. l + 26 = 17 C. 17l = 26 D. l + 17 = 26

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

I + 17 = 26

OpenStudy (steve816):

yeah, that's totally the answer.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Find it by doing you know that 26 - 17 will give you the number of games, so you set 26 - 17 = i, then set it equal to 26 by adding 17 to both sides to get i + 17 = 26, got it?